# Some models are useful, but for how long?: A decision theoretic approach to choosing when to refit large-scale prediction models

### v1

@@ 2.2 Risk and Ambiguity Aversion
This gives a simple approach for choosing $\lambda$ and $\theta$ that controls the Type I error rate under a null hypothesis that there is no drift. Alternative approaches based on likelihood-ratio testing [9], robust control theory [7], and robust sequential hypothesis testing [37] could also be used.
@@ 2.3 Estimating the Mean Squared Errors of Each Strategy
We evaluate the mean squared errors for each of the three model maintenance strategies when conducting downstream inference using predicted outcomes from our upstream AI/ML model. That is, we use our AI/ML model to generate predicted outcomes, $\hat{Y}$, which are then carried forward to a model that estimates a parameter, $\beta$, describing the relationship between a set of features, $X$, and $\hat{Y}$. In a randomized experiment, for example, $\beta$ is the effect of the treatment. We will denote the objective function for this regression as
$\hat{\beta}=\arg\min_{\beta}g\left(\hat{Y},X;\beta\right).$
Let $g$ be a generalized linear model (e.g., $g=||\hat{Y}-X\beta||_{2}^{2}$ in linear regression), however, this approach can be adapted to any model with a convex objective function with a similar strategy. Let $\hat{Y}_{ret}$, $\hat{Y}_{ref}$, $\hat{Y}_{rec}$, and $Y$ are the predicted $Y$ under the retain, refit, and recalibrate strategies, and true $Y$, respectively. Further, let $\hat{\beta}_{ret}$, $\hat{\beta}_{ref}$, $\hat{\beta}_{rec}$, and ${\beta}$ represent the parameter vectors.
Since the MSE depends on the number of labeled/unlabeled observations we collect, we also need to specify the relationship between budget and data collection. Let our total budget be $c$, let model refitting cost be $c_{model}$, let collecting $(X,Y)$ together cost $c_{labeled}$, and let collecting $X$ alone cost $c_{unlabeled}$. If one retains a model, no model refitting cost is needed, and one can use the entire budget to collect $\frac{c}{c_{unlabeled}}$ samples to estimate $\beta$. If one refits a model, one must first spend $c_{model}$ for refitting, leaving a budget of $\frac{c-c_{model}}{c_{unlabeled}}$ samples. Finally, if one chooses to recalibrate, choose the optimal ratio of labeled to unlabeled data ( costing $\frac{c}{c_{labeled}}$ and $\frac{c}{c_{unlabeled}}$) respectively to best estimate $\beta$ (we specify how to get this ratio below).
@@ 2.3.1 Strategy 1: Retain the Model
The simplest strategy is to continue using the original AI/ML model. In such a case, we would expect the difference between $\beta$ and $\hat{\beta}_{ret}$ to grow with time as the model drifts. To measure the rate at which this difference grows, we shall assume that post-initial training, the maintainers have collected data at regular time intervals to evaluate the performance of the model post-deployment. This is referred to as the monitoring, maintenance, verification, or, the moniker we shall use, the calibration phase [38, 21]. During the calibration phase we shall repeatedly compare $\beta$ and $\hat{\beta}_{ret}$ and fit a time series model to estimate the expected MSE at $t_{future}$. This allows us to extrapolate out how well retaining the model works from the calibration data, where we observe data, to $t_{future}$ where we are yet to observe data and a decision needs to be made. This can be done via the algorithm below:
@@ 2. Compute $\hat{\beta}_{ret,t}=\arg\min_{\beta}g(Y_{t},X_{t};\beta)$
The bootstrap sample size of $c/c_{unlab}$ illustrates the fact that since retaining does not require model refitting or additional labeled data, we get our best estimate for the MSE by spending it all on the unlabeled data, pushing it through our retained model, and computing its MSE.
@@ 2.3.2 Strategy 2: Refit the Model
If AI/ML model performance has degraded considerably, the best option may be to retire the old model and refit a new one. We make the assumption that any newly fit AI model, when trained on the new data, will have similar accuracy to the previous model. To refit, one must spend $c_{model}$ of their budget to train a new model and the remaining $(c-c_{model})/c_{unlab}$ for unlabeled samples to estimate the accuracy of the estimates using the new model. Since $g$ is a member of the generalized linear model family, we have, via Agresti [2], that
@@ 2.3.3 Strategy 3: Recalibrate the Model
The middle ground between retaining and refitting is to recalibrate the model. This requires us to first spend our budget on collecting an appropriate ratio of unlabeled and labeled data. Regressing the AI/ML-estimated outcomes from the unlabeled data against X yields a biased estimate of the parameter of interest. Using the labeled data, we compare our AI/ML generated $\hat{Y}$ to real $Y$ to learn the appropriate correction factor. Combining the biased parameter estimate with its appropriate correction factor to conduct valid statistical inference is what [20] call “Inference on Predicted Data.” Here, we utilize a mature procedure for such inference known as “PPI++” Angelopoulos et al. [5, 6] but other approaches that yield valid parameter estimates will work as well. Given the associated loss function $l(X,Y;\beta)$ for estimator of $\beta$, PPI++ is a parameter estimation procedure that minimizes:
@@
where $\gamma\in[0,1]$ is a data-estimated tuning parameter. When $\gamma$ is close to 0, this indicates that the AI model is perfectly accurate and the parameter estimates require no correction. Using the consistency results in Angelopoulos et al. [6] allow us to decompose the MSE via:
@@
where $(X_{lab}^{T}\hat{D}_{ref}X_{lab})$ is the information matrix for $\hat{\Delta}$. To extend this recalibration procedure to our temporal setting, we make the following extension to the PPI++ procedure:
AI/ML Model Recalibration Algorithm
@@
    Using said time series, estimate the forested distribution for for $\hat{W}_{i,refit},\hat{\Delta}_{i,ref},$ and $\hat{\gamma}_{i}$ at $t_{future}$.
@@
    For each sample size ratios $\zeta\in(0,...,1)$:
@@ 3 Simulation Study
We start by showing a simple simulation in which we know the ground truth. We generate 10,000 samples of $X_{1},X_{2}$, independent standard normal variables. We then generate
@@
where $I(\cdot)$ is the indicator function, and $\epsilon\sim N(0,0.05)$. We start with a model that is trained to predict $Y$ from $X_{1}$ and $X_{2}$ at time $t=0$. Then a downstream model estimates the relationship between $X_{1}$ and $Y$ via a linear regression, $\hat{Y}=X_{1}\hat{\beta}+\hat{\epsilon}$.
If the prediction model is accurate at time $t$, $\hat{\beta}$ should be close to $\exp(0.25t)$. As $t$ increases, the prediction model will be less accurate, and our model evaluation strategy will change from recalibrate to refit. As our data generating mechanism contains a step function, a reasonable choice for the prediction model is to use random forests trained with 5-fold cross-validation.
We simulate a calibration period between times $t=0$ and $t=3$, where we have 50 distinct points in which 200 labeled and unlabeled samples were collected. We seek to make a decision to refit, recalibrate, or retain at time point $t=3.5$. Our total budget to make a decision is \$200, the cost of model fitting was \$10, unlabeled data, $(X_{1,unlab},X_{2,unlab})$, each cost \$1 to collect, and labeled data, $(X_{1,lab},X_{2,lab},Y_{lab})$, cost \$2. As we are interested in a parameter from linear regression, $\hat{W}$ and $\hat{D}_{refit}$ has a simple representation
@@
and $X=[X_{1},X_{2}]$. The approximation here comes from how the classical $\hat{\sigma}_{unlab}(X_{unlab}^{T}X_{unlab})^{-1}$ estimator for the variance-covariance matrix is known to be biased when $Y$ is nonlinear. The common solution to this is to employ the sandwich estimator [51, 50, 25], but in practice, we find the extra number of parameters this requires makes it ill-suited for the extrapolation we require.
@@
With the settings above, each MSE estimation algorithm was run with 100 bootstrap samples on an M2 Max Macbook Pro with 32 GB of RAM, giving us the mean MSEs and variances as seen in Table 1. From this, we see that the retain strategy has the worst average MSE, followed by recalibrate and then refit. However, the recalibrate MSE has lower variance than the refit MSE, so our preference between the two is determined by how averse we are to ambiguity. In Figure 1, we evaluate our decision rules with $\lambda$ varying from 0 to 0.00125 and $\theta$ varying from 0 to 0.0012. We find a clear trade-off where one prefers to recalibrate when uncertainty aversion is high and refit when ambiguity aversion is high. The utility function chosen by our Type I error rate control with $\alpha=0.05$ indicates a preference for recalibration (Figure 1, left panel, black cross). At no point is retention the optimal choice under these simulation settings.
Figure 1: (Left): Our estimate of the best decision for various levels of aversion to uncertainty ($\theta$) and ambiguity ($\lambda$). The hypothesis-driven choice for $\lambda$, $\theta$ in Section 2.2 is marked with a square. (Middle): $\frac{\hat{w}_{ref}}{w_{rec}}$ and $\lambda$ for various levels of $\theta$. Values above 1 choose refit as optiomal. Above the horizontal line denotes refit, below is recalibrate. (Right): $\frac{\hat{w}_{ref}}{w_{rec}}$ for various $\theta$’s and $\lambda$’s. Above the horizontal line is refit, below is recalibrate.
@@ 4 Electricity Forecasting in New South Wales
As an example of decision-making using this procedure, we put ourselves in the situation of a large data center in New South Wales, Australia that requires 2000MWh of electricity [1]. In early 2024, this is a little over half of the capacity of the largest data center in the world [17]. To estimate electrical costs, the company created a two-step model forecasting electrical costs as follows:
  1. 1.
    Step 1: Demand Forecasting Every half hour, our AI/ML demand forecasting model estimates electrical demand $\hat{D}$ using data from the previous 7 days.
  2. 2.
    Step 2: Elasticity Estimation All forecasts for the day are regressed against price to obtain and estimate the daily price elasticity, $\beta$ from the equation: $\hat{D}=\beta X_{price}+\hat{\epsilon}.$
All estimation will be done using the Elec2 dataset [19] downloaded from Kaggle [39]. The data were pre-scaled so traditional log-log transformation was omitted [54]. The reliability of this procedure is based on the AI/ML demand model providing accurate forecasts of electrical demand. This makes this procedure sensitive to unpredictable shocks such as natural disasters or system failures. As an example of such a shock, midway through 1997, the electrical grid of New South Wales joined the grid in the neighboring state of Victoria to allow for load sharing. The decision maker, of course, may not always know when a shock (or gradual drift) occurs. We use a calibration period of 180 days with 60 days before grid unification and 120 days after. As we see the ground truth, we expect recalibrating to be appealing at first and then refitting to be preferable after unification. The decision-making process does not rely on knowing when, or if, unification has occurred a priori.
Our demand forecasting model is a long short-term memory model (LSTM) of 336 cells (equivalent to one week of demand data measured every 30 minutes) implemented in using the Keras machine learning library [13]. The model was trained on 11,150 datapoints at the start of the study. The LSTM had a $\tanh$ activation function, a sigmoid function for the recurrent activation function, and no dropout. All other parameters were set to the default as determined by the keras3 package in R [23].
Figure 2: (Left): Electrical price elasticity over time. (Right): MSE of estimated electrical price elasticity using the pre-unification forecast model. Vertical purple line represents the grid unification day. The elasticity and our ability to estimate rapidly decay, signifying a case of concept drift.
@@
Table 2: True values for mean MSE of $\beta$, estimated electrical bill for one month running the data center, and sum of differences between what we expected to pay versus what we paid

### v2

@@ 2.2.1 Choosing $\lambda$ and $\theta$
This gives a simple approach for choosing $\lambda$ and $\theta$ that controls the Type I error rate under a null hypothesis that there is no drift. Alternative approaches based on likelihood-ratio testing (Bonhomme and Weidner, 2022), robust control theory (Baltas, 2024), and robust sequential hypothesis testing (Quang, 1985) could also be used.
@@ 2.3 Estimating the Mean Squared Errors of Each Strategy
To employ the above decision rule, one needs a procedure to estimate the amount of risk and ambiguity that arises when estimating $\beta$ across each of the three strategies. To do this, we shall say that the parameter of interest $\beta$ comes from the minimization of the convex objective function:
$\hat{\beta}=\arg\min_{\beta}h\left(\hat{Y},X;\beta\right).$
In the succeeding sections, we shall assume that $h$ is the objective for a generalized linear model, however much of the intuition will hold for other convex estimation problems. Let $\hat{Y}_{ret}$, $\hat{Y}_{ref}$, $\hat{Y}_{rec}$, and $Y$ be the predicted $Y$ under the retain, refit, and recalibrate strategies, and true $Y$, respectively. Further, let $\hat{\beta}_{ret}$, $\hat{\beta}_{ref}$, $\hat{\beta}_{rec}$, and ${\beta}$ be the resulting parameter estimating resulting from the 3 strategies as well as the true $\beta$.
Since the MSE depends on the number of labeled/unlabeled observations we collect, we also need to specify the relationship between budget and data collection. Let our total budget be $c$, let model refitting cost be $c_{model}$, let collecting $(X,Y)$ together cost $c_{labeled}$, and let collecting $X$ alone cost $c_{unlabeled}$. If one retains a model, no model refitting cost is needed, and one can use the entire budget to collect $\frac{c}{c_{unlabeled}}$ samples to estimate $\beta$. If one refits a model, one must first spend $c_{model}$ for refitting, leaving a budget of $\frac{c-c_{model}}{c_{unlabeled}}$ samples. Finally, if one chooses to recalibrate, choose the optimal ratio of labeled to unlabeled data ( costing $\frac{c}{c_{labeled}}$ and $\frac{c}{c_{unlabeled}}$) respectively to best estimate $\beta$.
@@ 2.3.1 Strategy 1: Retain the Model
The simplest strategy is to use the original AI/ML model for everything. In such a case, we would expect the bias of $\hat{\beta}_{ret}$ to grow with increasing distance from the initial model fit (e.g. time, space, or distance). To measure the rate at which this bias grows, we shall assume that post-initial training, the maintainers have collected data at regular time intervals to evaluate the performance of the model post-deployment. This is referred to as the monitoring, maintenance, verification, or, the moniker we shall use, the calibration phase. During the calibration phase, we shall repeatedly compare $\beta$ and $\hat{\beta}_{ret}$ and fit a time series model to estimate the expected MSE at $t_{future}$. This allows us to extrapolate out how well retaining the model works from the calibration data, where we observe data, to $t_{future}$ where we are yet to observe data and a decision needs to be made. This can be done via the algorithm below:
@@ 2. Compute $\hat{\beta}_{ret,t}=\arg\min_{\beta}g(Y_{t},X_{t};\beta)$
The bootstrap sample size of $c/c_{unlab}$ illustrates the fact that since retaining does not require model refitting or additional labeled data, we get our best estimate for the MSE by spending it all on the unlabeled data, $X$, predicting $Y$ using our retained AI/ML model, and computing its MSE.
@@ 2.3.2 Strategy 2: Refit the Model
If AI/ML model performance has degraded considerably, the best option may be to retire the old model and refit a new one. We make the assumption that any newly fit AI/ML model, when trained on the new data, will have similar accuracy to the previous model. To refit, one must spend $c_{model}$ of their budget to train a new model and the remaining $(c-c_{model})/c_{unlab}$ for unlabeled samples to estimate the accuracy of the estimates using the new model. If we allow $h$ to be a generalized linear model, we have the traditional decomposition Agresti (2015):
@@ 2.3.3 Strategy 3: Recalibrate the Model
The middle ground between retaining and refitting is to recalibrate the model. This requires us to first spend our budget on collecting an appropriate ratio of unlabeled and labeled data. Regressing the AI/ML-predicted outcomes from the unlabeled data against $X$ yields a biased estimate of the parameter of interest. Using the labeled data, we compare our AI/ML generated $\hat{Y}$ to real $Y$ to learn the appropriate correction factor. Combining the biased parameter estimate with its appropriate correction factor to conduct valid statistical inference is what Hoffman et al. (2024) call “Inference on Predicted Data.” Here, we utilize a mature procedure for such inference known as “PPI++(Angelopoulos et al., 2023; Angelopoulos et al., 2024), but other approaches that yield valid parameter estimates will work as well. Given the associated loss function $l(X,Y;\beta)$ for estimator of $\beta$, PPI++ is a parameter estimation procedure that minimizes:
@@
where $\gamma\in[0,1]$ is a data-estimated tuning parameter and $\hat{\Delta}$ is the "Rectifier" as defined in Angelopoulos et al. (2024). When $\gamma$ is close to 0, this indicates that the AI model perfectly captures the Y|X relationship, and the parameter estimates require no correction. Using the unbiasedness of the Prediction-Powered Inference estimator (Angelopoulos et al., 2024), we can decompose the MSE via:
@@
where $(X_{lab}^{T}\hat{D}_{ref}X_{lab})$ is the information matrix for $\hat{\Delta}$. Note that $Var(\hat{\beta}_{ref})$ describes the amount of risk in the Prediction-Powered Inference estimate of $\beta$ while $Var(\hat{\Delta}_{ref})$ describes the amount of ambiguity. To extend this recalibration procedure to our temporal setting, we make the following extension to the PPI++ procedure:
Recalibrate Algorithm
@@
    Using said time series, estimate the predicted distribution for $\hat{W}_{i,refit},\hat{\Delta}_{i,ref},$ and $\hat{\gamma}_{i}$ at $t_{future}$.
@@
    For each sample size ratio $\zeta\in(0,...,1)$:
@@ 3 Simulation Study
We start by showing a simple simulation in which the ground truth is known. We generate 10,000 samples of $X_{1},X_{2}$, independent standard normal variables. We then generate
@@
where $I(\cdot)$ is the indicator function, and $\epsilon\sim N(0,0.05)$. We start with a model that is trained to predict $Y$ from $X_{1}$ and $X_{2}$ at time $t=0$. Then a downstream model estimates the relationship between $X_{1}$ and $Y$ via a linear regression, $\hat{Y}=X_{1}\hat{\beta}+\hat{\epsilon}$. If the prediction model is accurate at time $t$ (or location/distance t), $\hat{\beta}$ should be close to $\exp(0.25t)$. As $t$ increases, the prediction model will be less accurate, and our model evaluation strategy will change from recalibrate to refit. As our data-generating mechanism contains a step function, a reasonable choice for the prediction model is to use random forests trained with 5-fold cross-validation.
We simulate a calibration period between times $t=0$ and $t=3$, where we have 50 distinct points in which 200 labeled and unlabeled samples were collected. We seek the decision to refit, recalibrate, or retain at time point $t=3.5$. Our total budget to make a decision is 200, the cost of model-fitting was 10, unlabeled data, $(X_{1,unlab},X_{2,unlab})$, each cost 1 to collect, and labeled data, $(X_{1,lab},X_{2,lab},Y_{lab})$, cost 2. As we are interested in a parameter from linear regression, $\hat{W}$ and $\hat{D}_{refit}$ has a simple representation:
@@
and $X=[X_{1},X_{2}]$. The approximation here comes from how the classical $\hat{\sigma}_{unlab}(X_{unlab}^{T}X_{unlab})^{-1}$ estimator for the variance-covariance matrix is known to be biased when $Y$ is nonlinear. The common solution to this is to employ the sandwich estimator (White, 1982, 1980; Liang and Zeger, 1986), but in practice, we find the extra number of parameters this requires makes it ill-suited for the extrapolation we require.
@@
With the settings above, each MSE estimation algorithm was run with 100 bootstrap samples on an M2 Max Macbook Pro with 32 GB of RAM, giving us the mean MSEs and variances as seen in Table 1. From this, we see that the retain strategy has the worst average MSE, followed by recalibrate and then refit. However, the recalibrate MSE has lower variance than the refits MSE, so our preference between the two is determined by how averse we are to ambiguity. In Figure 1, we evaluate our decision rules with $\lambda$ varying from 0 to 0.00125 and $\theta$ varying from 0 to 0.0012. We find a clear trade-off where one prefers to recalibrate when uncertainty aversion is high and refit when ambiguity aversion is high. The utility function chosen by our Type I error rate control with $\alpha=0.05$ indicates a preference for recalibration (Figure 1, left panel, black cross). At no point is retention the optimal choice under these simulation settings.
Figure 1: (Left): Our estimate of the best decision for various levels of aversion to uncertainty ($\theta$) and ambiguity ($\lambda$). The hypothesis-driven choice for $\lambda$, $\theta$ in Section 2.2.1 is marked with a square. (Middle): $\frac{\hat{w}_{ref}}{w_{rec}}$ and $\lambda$ for various levels of $\theta$. Values above 1 choose refit as optimal. Above the horizontal line denotes refit, below is recalibrate. (Right): $\frac{\hat{w}_{ref}}{w_{rec}}$ for various $\theta$’s and $\lambda$’s. Above the horizontal line is refit, below is recalibrate.
@@ 4 Electricity Forecasting in New South Wales
As an example of decision-making using this procedure, we put ourselves in the situation of a large data center in New South Wales, Australia that requires 2000MWh of electricity (AFL, 2023). In early 2024, this is a little over half of the capacity of the largest data center in the world  (Gooding, 2024). To estimate electrical costs, the company created a two-step model forecasting electrical costs. First, every half hour, demand was forecasted for electrical demand $\hat{e}$ using data from the previous 7 days. Second, forecasts for the day are regressed against price to obtain and estimate the daily price elasticity, $\beta$ from the equation: $\hat{e}=\beta X_{price}+\hat{\epsilon}.$All estimation was done using the Elec2 dataset (Harries, 1995) downloaded from Kaggle (Sharan, 2020). The data were pre-scaled so traditional log-log transformation was omitted (Wooldridge, 2010). The reliability of this procedure is based on the AI/ML demand model, which provides accurate forecasts of electrical demand. This makes this procedure sensitive to unpredictable shocks such as natural disasters or system failures. As an example of such a shock, midway through 1997, the electrical grid of New South Wales joined the grid in the neighboring state of Victoria to allow for load sharing. Grid unification leads to changes in both the total supply and total demand so is an example of concept drift changing the relationship between the elasticity (the outcome) and the time of day (the covariate of interest). The decision maker, of course, may not always know when a shock (or gradual drift) occurs. We use a calibration period of 180 days with 60 days before grid unification and 120 days after. As we see the ground truth, we expect recalibrating to be appealing at first and then refitting to be preferable after unification. The decision-making process does not rely on knowing when, or if, unification has occurred a priori.
Our demand forecasting model is a long short-term memory model (LSTM) of 336 cells (equivalent to one week of demand data measured every 30 minutes) implemented using the Keras machine learning library. The model was trained on 11,150 data points at the start of the study. The LSTM had a $\tanh$ activation function, a sigmoid function for the recurrent activation function, and no dropout. All other parameters were set to the default as determined by the keras3 package in R.
Figure 2: (Left): Electrical price elasticity over time. (Right): MSE of estimated electrical price elasticity using the pre-unification forecast model. The vertical purple line represents the grid unification day. The elasticity and our ability to estimate rapidly decay, signifying a case of concept drift.
@@
Table 2: True values for mean MSE of $\beta$, estimated electrical bill for one month running the data center, and the sum of differences between what we expected to pay versus what we paid
@@
5 Google Flu Trends
Google Flu Trends was a high-profile initiative launched in 2008 that used large-scale tracking of Google search queries about illness-related terms as a low-cost way to estimate the prevalence of influenza-like illness in the United States (Ginsberg et al., 2009). It initially launched to great excitement due to its impressive predictive accuracy during the 2003-2008 flu seasons as well as its cheaper costs compared to the existing national-wide flu monitoring system, the Center for Disease Control and Prevention’s Influenza Hospitalization Surveillance Network. However, after a series of high-profile failures, such as overestimation of twice the true number of flu-like illnesses during the 2012 and 2013 season (Lazer et al., 2014), and demonstrations that Google’s predictions did not consistently outperform a simple forward projection of 2-week-old CDC data (Katsikopoulos et al., 2022), this lead to the project’s abandonment in 2015.
While many analyses have been done on the project’s failure, a consistent theme was how static Google’s forecasting model was (Lazer et al., 2014; Butler, 2013; Lazer, 2015; Franzén, 2023). Unlike models whose parameters can dynamically change over time as the diseases ebb and flow, the Google Flu Trends model was fit on weekly CDC data from 2003-2008 and minus an ill-documented update in 2009, was never refit or retrained till abandonment. This made the model ill-suited to new events such as the off-season 2009 spring epidemic (Schmidt, 2019), media-stoaked panic about the flu during the 2013 season (Lazer et al., 2014), and changes in Google’s search algorithm (of which the official Google search blog reported 86 changes in their search algorithm from June and July 2012 alone (Lazer et al., 2014)). This makes this a mix of both gradual and fast concept drifts as the relationship between the ILI counts and the covariates of epidemiological interest (in this case geographic region).
For this reason, Google Flu Trends serves as a good example of the importance of making sure an algorithm is kept up-to-date. To this end, we would like to recreate the problem these Google Researchers face: “How often should we be refitting the Google Flu model to new data?” and demonstrate how our procedure could have been used to suggest a refitting time before the catastrophically wrong predictions during the 2012-2013 seasons occurred.
We obtained weekly data of the CDC and Google estimated number of influenza-like illness (ILI) cases from 2003-2015 for each of the 10 Health and Human Services Regions using the epidatr R package from the Delphi project (McDonald and DeFries, 2024). While changes in methodology since 2015 meant that the CDC’s numbers now differ substantially from Google’s past estimates, we found that taking a region-specific linear transformation of Google’s estimates gave us a correlation between predicted and true ILI counts that matched those of the original paper (range: 0.79-0.9 vs 0.80-0.96). The data was then divided into three time periods to mirror the information the Google Flu team had access to. Data from 2003-2006 were treated as training data. We considered 3 different lengths of calibration periods 1) January 1, 2006– June 30, 2008 (soon after the training period), 2) January 1, 2006 – June 30, 2009 (in between), and 3) January 1, 2008 – June 30, 2011 (right before big misses of 2012-2013). Budget costs for this procedure are difficult to determine as Google did not release cost estimates for their model, so to be conservative, we will have CDC data only cost twice as much as Google data to produce. We will thus set a total budget of 200, the model training cost was set to 10, and unlabeled and labeled data were given costs of 1 and 2 respectively. The upstream, black-box model was Google Flu Trend’s ILI estimates with our region-specific recalibration. The regression of scientific interest is ILI regressed on region, specifically the regression coefficient for Region 2 which covers New York, New Jersey, Puerto Rico, and the Virgin Islands. Thus, each procedure evaluates if one is interested in obtaining the best estimates of ILI counts for the upcoming flu season in Region 2, one should retain, refit, or recalibrate their model.
Figure 4: Decision space for the Refit vs Recalibrate for Google Flu Trends in 2008 (left) 2009 (middle) and 2011 (right) across different uncertainty and ambiguity aversion parameters.
Figure 4 shows the decision space for our analysis of the 2008, 2009, and 2011 flu seasons. In 2008 we can see that the decision space is almost completely in favor of keeping the model with only those with no aversion to ambiguity favoring a refit. In 2009, we now see an equal division of the space between those who favor refit versus those who believe a recalibration is a better approach. This increase is indicative of the model getting more and more out of date. Finally, we see by 2011, only one year before the major failures of the 2012-2013 season, almost all the results were in favor of refitting the model. Taken together, our analysis corroborates the existing findings that a decrease in model accuracy over time played a major role in Google Flu Trends missing the 2011-2012 seasons. While determining which exact aversion to uncertainty and ambiguity would work best for Google requires knowledge of Google’s risk preferences that we are not privacy to, across a large range of preferences, it seems that Google Flu Trends would have been best served by refitting their model sometime between 2008-2009 and definitely before 2011. Moreover, this framework was not only able to give an explicit choice between strategies for maintaining the Google Flu model but was able to make this decision using the information that was available at the time, not as a part of a retrospective assessment.